# Monadic reconstruction of unitary Drinfeld centers and Factorization Homology

### v1

@@ 3.3. Monadic reconstruction of the algebraic Drinfeld center
$\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega\in\operatorname{ONB}(b,a\odot U)}(\mathrm{id}_{U}\odot\omega^{\vee}\odot\mathrm{id}_{X}\odot\omega^{*})(\bar{R}_{U}\odot\mathrm{id}_{\bar{a}}\odot\mathrm{id}_{X}\odot\mathrm{id}_{a})(\eta_{X}\odot\mathrm{id}_{U})\ ,$
@@
$\sum_{\omega\in\operatorname{ONB}(b,U)}(\mathrm{id}_{U}\odot\omega^{\vee}\odot\mathrm{id}_{X}\odot\omega^{*})(\bar{R}_{U}\odot\mathrm{id}_{X})\ .$
@@ 3.4. The adjunction
Given $f\in\operatorname{Hom}_{\operatorname{Vec}(\mathcal{C})}(X,X^{\prime})$ and $g\in\operatorname{Hom}_{\operatorname{Vec}(\mathcal{C})}(Y,Y^{\prime})$, $Z(f)\in\operatorname{Hom}_{\mathcal{Z}\operatorname{Vec}(\mathcal{C})}(Z(X),Z(X^{\prime}))$, and since $\beta^{X}$ is a half-braiding, $\beta^{X}_{Y^{\prime}}(\mathrm{id}_{Z(X)}\odot g)=(g\odot\mathrm{id}_{Z(X)})\beta^{X}_{Y}$. Therefore
@@
$\varepsilon_{(M,\gamma)}\in\operatorname{Hom}_{\mathcal{Z}\operatorname{Vec}(\mathcal{C})}((Z_{\operatorname{Vec}}(M),\beta^{M}),(M,\gamma))\ .$
@@
We shall show that, for any $X\in\operatorname{Vec}(\mathcal{C})$ and any morphism $f\in\operatorname{Hom}_{\mathcal{Z}\operatorname{Vec}(\mathcal{C})}((M,\gamma),(N,\gamma^{\prime}))$,
@@
By Lemma 3.13, this will show $\varepsilon_{(N,\gamma^{\prime})}Z(f)=f\varepsilon_{(M,\gamma)}$ and therefore the naturality of $\varepsilon$.
@@
For all $X\in\operatorname{Vec}(\mathcal{C})$ we have $\varepsilon_{\mathcal{F}(X)}Z_{\operatorname{Vec}}(\eta_{X})=\mathrm{id}_{\mathcal{F}(X)}$, where $\varepsilon_{\mathcal{F}(X)}=\varepsilon_{(Z(X),\beta^{X})}$.
@@
We shall show that $(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)}Z_{\operatorname{Vec}}(\eta_{X}))\partial_{X,U}=\partial_{X,U}$ for each $U\in\mathcal{C}$ and invoke Lemma 3.13 to conclude the proof. We have
$(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)}Z_{\operatorname{Vec}}(\eta_{X}))\partial_{X,U}$ $=(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)}Z_{\operatorname{Vec}}(\eta_{X}))\beta^{X}_{U}(\eta_{X}\odot\mathrm{id}_{U})$
$=(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)})\beta^{Z_{\operatorname{Vec}}(X)}_{U}(\eta_{Z_{\operatorname{Vec}}(X)}\eta_{X}\odot\mathrm{id}_{U})$
$=(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)})\partial_{Z_{\operatorname{Vec}}(X),U}(\eta_{X}\odot\mathrm{id}_{U})$
@@
$\mu_{X}=\sum_{a,b,c\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega\in\operatorname{ONB}(c,a\odot b)}(\omega^{\vee}\odot\mathrm{id}_{X}\odot\omega^{*})\ ,$
@@ 3.5. Linear isomorphism
There is a functor $\Phi:\mathcal{Z}\operatorname{Vec}(\mathcal{C})\to Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}$
@@
There is a functor $Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}\to\mathcal{Z}\operatorname{Vec}(\mathcal{C})$.
@@
is an inverse for $\beta^{\tau}_{U}$. Since, by naturality, $\beta^{\tau}$ is determined by the family $\{\beta^{\tau}_{U}\}_{U\in\mathcal{C}}$, and inverse of $\beta^{\tau}$ is also defined by its values at objects $U\in\mathcal{C}$. Thus, $\tilde{\beta}=\{\tilde{\beta}^{\tau}_{U}\}_{U\in\mathcal{C}}$ has a unique extension to a natural transformation $(-)\odot M\to M\odot(-)$ between endofunctors of $\operatorname{Vec}(\mathcal{C})$, and this extension is $(\beta^{\tau})^{-1}$.
@@
Hence, defining $\Psi(M,\tau):=(M,\beta^{\tau})$ on object $(M,\tau)\in Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}$, and $\Psi(f)=f$ on morphisms, we get a functor $\Psi:Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}\to\mathcal{Z}\operatorname{Vec}(\mathcal{C})$. ∎
@@ 3.6. Bimonadic structure of $Z$
Now we shall use lemmas 3.13, 3.15 and 3.16 to give to $Z_{\operatorname{Vec}}$ the structure of a bimonad.
Let $X_{1},X_{2}\in\operatorname{Vec}(\mathcal{C})$. The collection
$\{(\partial_{X_{1},K}\odot\mathrm{id}_{Z_{\operatorname{Vec}}(X_{2})})(\mathrm{id}_{X_{1}}\odot\partial_{X_{2},K}):X_{1}\odot X_{2}\odot K\to K\odot Z_{\operatorname{Vec}}(X_{1})\odot Z_{\operatorname{Vec}}(X_{2})\}_{K\in\operatorname{Vec}(\mathcal{C})}$
is natural in $K\in\operatorname{Vec}(\mathcal{C})$. By Lemma 3.13, there is a unique $Z_{2}(X_{1},X_{2}):Z_{\operatorname{Vec}}(X_{1}\odot X_{2})\to Z_{\operatorname{Vec}}(X_{1})\odot Z_{\operatorname{Vec}}(X_{2})$ such that
$(\partial_{X_{1},U}\odot\mathrm{id}_{Z_{\operatorname{Vec}}(X_{2})})(\mathrm{id}_{X_{1}}\odot\partial_{X_{2},U})=(\mathrm{id}_{U}\odot Z_{2}(X_{1},X_{2}))\partial_{X_{1}\odot X_{2},U}$
@@
$Z_{\operatorname{Vec}}(X_{1}\odot X_{2})(c)=\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}\operatorname{Vec}(\mathcal{C})(c,\bar{a}\odot X_{1}\odot X_{2}\odot a),$
and therefore for each $v\in Z_{\operatorname{Vec}}(X_{1}\odot X_{2})(c)$ there are only finitely many $a\in\operatorname{Irr}(\mathcal{C})$ for which the image of $v$ under
@@
is non-zero.
By naturarility, it suffices to prove
$(\partial_{X_{1},U}\odot\mathrm{id}_{Z_{\operatorname{Vec}}(X_{2})})(\mathrm{id}_{X_{1}}\odot\partial_{X_{2},U})=(\mathrm{id}_{U}\odot Z_{2}(X_{1},X_{2}))\partial_{X_{1}\odot X_{2},U}$
@@
$(\partial_{X_{1},U}\odot\mathrm{id}_{Z_{\operatorname{Vec}}(X_{2})})(\mathrm{id}_{X_{1}}\odot\partial_{X_{2},U})$ $=\bar{R}_{c}\odot\mathrm{id}_{X_{1}}\odot\bar{R}_{c}\odot\mathrm{id}_{X_{2}}\odot c$
@@
The unique morphism $Z_{0}:Z_{\operatorname{Vec}}({\mathds{1}})\to{\mathds{1}}$ saitsfying
@@
Given $c\in\operatorname{Irr}(\mathcal{C})$ and $v\in Z_{\operatorname{Vec}}({\mathds{1}})(c)\simeq\bigoplus_{a}\mathcal{C}(c,\bar{a}\odot a)$, there are only finitely many $a\in\operatorname{Irr}(\mathcal{C})$ for which the image of $v$ under $R_{a}^{*}:\mathcal{C}(c,\bar{a}\odot a)\to\mathcal{C}(c,{\mathds{1}})$ is non-zero. Hence $Z_{0}$ is well-defined.
By naturality, it suffices to show that
@@
The pair $(Z_{2},Z_{0})$ is an op-lax structure on $Z_{\operatorname{Vec}}$, turning $(Z_{\operatorname{Vec}},\mu,\eta,Z_{2},Z_{0})$ into a bimonad in $\operatorname{Vec}(\mathcal{C})$.
By making use of the bimonadic structure on $Z_{\operatorname{Vec}}$ given by $(Z_{2},Z_{0})$, the category $Z_{\operatorname{Vec}}{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}$ can be given a tensor structure such that the forgetful functor $Z_{\operatorname{Vec}}{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}\to\operatorname{Vec}(\mathcal{C})$ is strictly monoidal. Indeed, if $(M,\tau)$ and $(N,\upsilon)$ are $Z_{\operatorname{Vec}}$-modules,
$Z_{\operatorname{Vec}}(M\odot N)\overset{Z_{2}}{\longrightarrow}Z_{\operatorname{Vec}}(M)\odot Z_{\operatorname{Vec}}(N)\overset{\tau\odot\upsilon}{\longrightarrow}M\odot N$
defines a $Z_{\operatorname{Vec}}$-module structure on $M\odot N$.
@@
Let $(M,\tau)$ and $(N,\upsilon)$ be $Z_{\operatorname{Vec}}$-modules. Then
@@
The linear isomorphism $Z_{\operatorname{Vec}}{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}\simeq\mathcal{Z}\operatorname{Vec}(\mathcal{C})$ is an isomorphism of tensor categories.
@@ 4.1. Bounded operators on Hilbert space objects
$\mathbb{B}^{l}_{H}(a):=\mathrm{Hilb}(\mathcal{C})(a\otimes H,H)\ \ \text{and}\ \ \mathbb{B}^{r}_{H}(a):=\mathrm{Hilb}(\mathcal{C})(H\otimes a,H).$
@@
$\mathcal{L}^{l}_{V}(-)=\operatorname{Vec}(\mathcal{C})((-)\odot V,V)\ .$
@@
The proof of the first claim is identitical to the proof of the second. We only explain the latter.
For simplicity, we shall ommit the notation $\operatorname{For}$ for the forgetful functor $\mathrm{Hilb}(\mathcal{C})\to\operatorname{Vec}(\mathcal{C})$. There is a canonical linear map
@@
Using the generic notation $\triangleright:\mathbb{B}^{l}_{H}\odot H\to H$ for the morphism defined above, and denoting by $\mu:\mathbb{B}^{l}_{H}\odot\mathbb{B}^{l}_{H}\to\mathbb{B}^{l}_{H}$ the multiplication morphism, we want to show that the diagram

### v2

@@ 3.3. Monadic reconstruction of the algebraic Drinfeld center
$\sum_{a\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega\in O(b,a\odot U)}(\mathrm{id}_{U}\odot\omega^{\vee}\odot\mathrm{id}_{X}\odot\omega^{*})(\bar{R}_{U}\odot\mathrm{id}_{\bar{a}}\odot\mathrm{id}_{X}\odot\mathrm{id}_{a})(\eta_{X}\odot\mathrm{id}_{U})\ ,$
@@
$\sum_{\omega\in O(b,U)}(\mathrm{id}_{U}\odot\omega^{\vee}\odot\mathrm{id}_{X}\odot\omega^{*})(\bar{R}_{U}\odot\mathrm{id}_{X})\ .$
@@ 3.4. The adjunction
Given $f\in\operatorname{Vec}(\mathcal{C})(X,X^{\prime})$ and $g\in\operatorname{Vec}(\mathcal{C})(Y,Y^{\prime})$, $Z(f)\in\mathcal{Z}\operatorname{Vec}(\mathcal{C})(Z(X),Z(X^{\prime}))$, and since $\beta^{X}$ is a half-braiding, $\beta^{X}_{Y^{\prime}}(\mathrm{id}_{Z(X)}\odot g)=(g\odot\mathrm{id}_{Z(X)})\beta^{X}_{Y}$. Therefore
@@
$\varepsilon_{(M,\gamma)}\in\mathcal{Z}\operatorname{Vec}(\mathcal{C})((Z(M),\beta^{M}),(M,\gamma))\ .$
@@
We shall show that, for any $X\in\operatorname{Vec}(\mathcal{C})$ and any morphism $f\in\mathcal{Z}\operatorname{Vec}(\mathcal{C})((M,\gamma),(N,\gamma^{\prime}))$,
@@
By Lemma 3.13, this implies $\varepsilon_{(N,\gamma^{\prime})}Z(f)=f\varepsilon_{(M,\gamma)}$ and therefore the naturality of $\varepsilon$.
@@
For all $X\in\operatorname{Vec}(\mathcal{C})$ we have $\varepsilon_{\mathcal{F}(X)}Z(\eta_{X})=\mathrm{id}_{\mathcal{F}(X)}$, where $\varepsilon_{\mathcal{F}(X)}=\varepsilon_{(Z(X),\beta^{X})}$.
@@
We shall show that $(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)}Z(\eta_{X}))\partial_{X,U}=\partial_{X,U}$ for each $U\in\mathcal{C}$ and invoke Lemma 3.13 to conclude the proof. We have
$(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)}Z(\eta_{X}))\partial_{X,U}$ $=(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)}Z(\eta_{X}))\beta^{X}_{U}(\eta_{X}\odot\mathrm{id}_{U})$
$=(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)})\beta^{Z(X)}_{U}(\eta_{Z(X)}\eta_{X}\odot\mathrm{id}_{U})$
$=(\mathrm{id}_{U}\odot\varepsilon_{\mathcal{F}(X)})\partial_{Z(X),U}(\eta_{X}\odot\mathrm{id}_{U})$
@@
$\mu_{X}=\sum_{a,b,c\in\operatorname{Irr}(\mathcal{C})}\sum_{\omega\in O(c,a\odot b)}(\omega^{\vee}\odot\mathrm{id}_{X}\odot\omega^{*})\ ,$
@@ 3.5. Linear isomorphism
There is a canonical functor $\Phi:\mathcal{Z}\operatorname{Vec}(\mathcal{C})\to Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}$
@@
There is a canonical functor $\Psi:Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}\to\mathcal{Z}\operatorname{Vec}(\mathcal{C})$.
@@
is an inverse for $\beta^{\tau}_{U}$. By naturality, $\beta^{\tau}$ is determined by the family $\{\beta^{\tau}_{U}\}_{U\in\mathcal{C}}$, and the inverse of $\beta^{\tau}$ is also defined by its values at objects $U\in\mathcal{C}$. Thus, $\tilde{\beta}=\{\tilde{\beta}^{\tau}_{U}\}_{U\in\mathcal{C}}$ has a unique extension to a natural transformation $(-)\odot M\to M\odot(-)$ between endofunctors of $\operatorname{Vec}(\mathcal{C})$, and this extension is $(\beta^{\tau})^{-1}$.
@@
Hence, defining $\Psi(M,\tau):=(M,\beta^{\tau})$ on objects $(M,\tau)\in Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}$, and $\Psi(f)=f$ on morphisms, we get a functor $\Psi:Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}\to\mathcal{Z}\operatorname{Vec}(\mathcal{C})$. ∎
@@ 3.6. Bimonadic structure of $Z$
Now we shall use lemmas 3.13, 3.15 and 3.16 to give to $Z$ the structure of a bimonad. Let $X_{1},X_{2}\in\operatorname{Vec}(\mathcal{C})$. The collection
$\{(\partial_{X_{1},K}\odot\mathrm{id}_{Z(X_{2})})(\mathrm{id}_{X_{1}}\odot\partial_{X_{2},K}):X_{1}\odot X_{2}\odot K\to K\odot Z(X_{1})\odot Z(X_{2})\}_{K\in\operatorname{Vec}(\mathcal{C})}$
is natural in $K\in\operatorname{Vec}(\mathcal{C})$. By Lemma 3.13, there is a unique $Z_{2}(X_{1},X_{2}):Z(X_{1}\odot X_{2})\to Z(X_{1})\odot Z(X_{2})$ such that
$(\partial_{X_{1},U}\odot\mathrm{id}_{Z(X_{2})})(\mathrm{id}_{X_{1}}\odot\partial_{X_{2},U})=(\mathrm{id}_{U}\odot Z_{2}(X_{1},X_{2}))\partial_{X_{1}\odot X_{2},U}$
@@
$Z(X_{1}\odot X_{2})(c)=\bigoplus_{a\in\operatorname{Irr}(\mathcal{C})}\operatorname{Vec}(\mathcal{C})(c,\bar{a}\odot X_{1}\odot X_{2}\odot a),$
and therefore for each $v\in Z(X_{1}\odot X_{2})(c)$ there are only finitely many $a\in\operatorname{Irr}(\mathcal{C})$ for which the image of $v$ under
@@
is non-zero. By naturarility, it suffices to prove
$(\partial_{X_{1},U}\odot\mathrm{id}_{Z(X_{2})})(\mathrm{id}_{X_{1}}\odot\partial_{X_{2},U})=(\mathrm{id}_{U}\odot Z_{2}(X_{1},X_{2}))\partial_{X_{1}\odot X_{2},U}$
@@
$(\partial_{X_{1},U}\odot\mathrm{id}_{Z(X_{2})})(\mathrm{id}_{X_{1}}\odot\partial_{X_{2},U})$ $=\bar{R}_{c}\odot\mathrm{id}_{X_{1}}\odot\bar{R}_{c}\odot\mathrm{id}_{X_{2}}\odot c$
@@
The unique morphism $Z_{0}:Z({\mathds{1}})\to{\mathds{1}}$ saitsfying
@@
Given $c\in\operatorname{Irr}(\mathcal{C})$ and $v\in Z({\mathds{1}})(c)\simeq\bigoplus_{a}\mathcal{C}(c,\bar{a}\odot a)$, there are only finitely many $a\in\operatorname{Irr}(\mathcal{C})$ for which the image of $v$ under $R_{a}^{*}:\mathcal{C}(c,\bar{a}\odot a)\to\mathcal{C}(c,{\mathds{1}})$ is non-zero. Hence $Z_{0}$ is well-defined. By naturality, it suffices to show that
@@
The pair $(Z_{2},Z_{0})$ is an op-lax structure on $Z$, turning $(Z,\mu,\eta,Z_{2},Z_{0})$ into a bimonad in $\operatorname{Vec}(\mathcal{C})$.
By making use of the bimonadic structure on $Z$ given by $(Z_{2},Z_{0})$, the category $Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}$ can be given a tensor structure such that the forgetful functor $Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}\to\operatorname{Vec}(\mathcal{C})$ is strictly monoidal. Indeed, if $(M,\tau)$ and $(N,\upsilon)$ are $Z$-modules,
$Z(M\odot N)\overset{Z_{2}}{\longrightarrow}Z(M)\odot Z(N)\overset{\tau\odot\upsilon}{\longrightarrow}M\odot N$
defines a $Z$-module structure on $M\odot N$.
@@
Let $(M,\tau)$ and $(N,\upsilon)$ be $Z$-modules. Then
@@
The linear isomorphism $Z{\hbox{-}}\operatorname{Mod}_{\operatorname{Vec}(\mathcal{C})}\simeq\mathcal{Z}\operatorname{Vec}(\mathcal{C})$ is an isomorphism of tensor categories.
@@ 4.1. Bounded operators on Hilbert space objects
$\mathbb{B}^{l}_{H}(a):=\mathrm{Hilb}(\mathcal{C})(a\otimes H,H)\ \ \text{and}\ \ \mathbb{B}^{r}_{H}(a):=\mathrm{Hilb}(\mathcal{C})(H\otimes a,H),$
with multiplications $\mu^{l}$ and $\mu^{r}$ given by
$\mu^{l}(f\odot g):=f\circ(\mathrm{id}_{a}\otimes g)\ \ ,\ \ \mu^{r}(f^{\prime}\odot g^{\prime}):=g\circ(f\otimes\mathrm{id}_{b})\ ,$
for $(f,g)\in\mathbb{B}^{l}_{H}(a)\times\mathbb{B}^{l}_{H}(b)$ and $(f^{\prime},g^{\prime})\in\mathbb{B}^{r}_{H}(a)\times\mathbb{B}^{r}_{H}(b)$, and *-structures given by
$j^{l}(f):=\Phi^{l}_{a}(f)^{*}\ \ ,\ \ j^{r}(f^{\prime}):=\Phi^{r}_{a}(f^{\prime})^{*}\ ,$
using the Frobenius reciprocity maps. That $\mathbb{B}^{l}_{H}$ and $\mathbb{B}^{r}_{H}$ are W-algebra objects is seen by realizing that they correspond to the cyclic right $\mathcal{C}$-module W-category and the cyclic left $\mathcal{C}$-module W-category generated by $H$, respectively.
@@
$\mathcal{L}^{l}_{V}(-)=\operatorname{Vec}(\mathcal{C})((-)\odot V,V)\ ,$
with algebra structure analogous to that of $\mathbb{B}^{l}_{H}$. The algebra object $\mathcal{L}^{r}_{V}$ is defined similarly.
@@
Recall that $\operatorname{For}$ denotes the forgetful functor $\mathrm{Hilb}(\mathcal{C})\to\operatorname{Vec}(\mathcal{C})$.
@@
The proof of the first claim is identitical to the proof of the second. We only explain the latter. For simplicity, we shall ommit the notation for the forgetful functor $\operatorname{For}:\mathrm{Hilb}(\mathcal{C})\to\operatorname{Vec}(\mathcal{C})$. There is a canonical linear map
@@
Using the generic notation $\triangleright:\mathbb{B}^{l}_{H}\odot H\to H$ for the morphism defined above, and denoting by $\mu:\mathbb{B}^{l}_{H}\odot\mathbb{B}^{l}_{H}\to\mathbb{B}^{l}_{H}$ the multiplication morphism for simplicity, we want to show that the diagram